In a certain group of computer personnel, 65% have insufficient knowledge of hardware, 45% have inadequate idea of software & 70% are in either one or both of the two categories. What is the percentage of people who do not know software among those who have a sufficient knowledge of hardware?

Answer by stanbon(75887)

In a certain group of computer personnel, 65% have insufficient knowledge of hardware, 45% have inadequate idea of software & 70% are in either one or both of the two categories. What is the percentage of people who do not know software among those who have a sufficient knowledge of hardware?
----
P(know soft |no hard) = P(soft AND hard)/P(hard) = 0.30/0.35 = 6/7
===========
Cheers,
Stan H.
==========